Test the following series for convergence,

Σ n xn-1 , x > 0


1
Expert's answer
2021-03-30T07:28:03-0400

Let us test the series "\sum_{n=1}^{\infty} n x^{n-1},\ x > 0," for convergence using D'Alembert criterion for positive series:


"\lim\limits_{n\to\infty}\frac{(n+1)x^n}{nx^{n-1}}=x\lim\limits_{n\to\infty}\frac{n+1}{n}=x\cdot 1=x"


If "0<x<1" then the series is convergent. If "x=1" then the series "\sum_{n=1}^{\infty} n" is divergent because "\lim\limits_{n\to\infty}n=\infty\ne 0."


Therefore, the series "\sum_{n=1}^{\infty} n x^{n-1},\ x > 0," is convergent if and only if "0<x<1".